OpenStudy (enricoanders):

Entomologists are biological scientists who study insects. Entomologists studying tree crickets have found that they chirp at different rates depending on the temperature. The number of chirps per minute, C, that the tree cricket makes is linearly dependent on the temperature, T, in Fahrenheit. The crickets do not chirp at all below 35 degrees and at 65 degrees they chirp about 138 times per minute. A) Express the number of chirps, C, as a function of the temperature, T. Be sure to use function notation:____________ B) How many chirps per minute will crickets make at 80 degrees?

OpenStudy (enricoanders):

@satellite73

satellite73 (satellite73):

find the equation of the line between the two points \((0,-35)\) and\((65,138)\) find the slope first, then the y intercept is given to you by the first point, so you don't need to find it

OpenStudy (enricoanders):

so the y intercept is gonna be -35 after i find the slope?

satellite73 (satellite73):

yes it is

OpenStudy (enricoanders):

so y=173/65x-35

satellite73 (satellite73):

it is going to be \[C=mT-35\] as soon as you find \(m\)

OpenStudy (enricoanders):

c=173/65t-35

satellite73 (satellite73):

hold on

OpenStudy (enricoanders):

what about part two? B) How many chirps per minute will crickets make at 80 degrees?

satellite73 (satellite73):

i may have made a mistake here

satellite73 (satellite73):

oops yes i did

satellite73 (satellite73):

the first is the temp, not the second the two points are \((-35,0)\) and \((65,138)\)

OpenStudy (enricoanders):

so c=23/5t-35

satellite73 (satellite73):

no the y intercept is not \(-35\) i screwed up

satellite73 (satellite73):

what did you get for the slope/

OpenStudy (enricoanders):

0 sorry

OpenStudy (enricoanders):

23/5

OpenStudy (enricoanders):

138/100=69/50

satellite73 (satellite73):

yeah that

OpenStudy (enricoanders):

c=69/50t-0

OpenStudy (enricoanders):

what about part 2 B) How many chirps per minute will crickets make at 80 degrees?

satellite73 (satellite73):

no you just got the slope you don't have the equation for the line yet you need the point slope formula

OpenStudy (enricoanders):

okay

satellite73 (satellite73):

\[C-0=\frac{69}{50}(x+35)\]

satellite73 (satellite73):

i would use decimals \[C=1.38(T+35)\]

OpenStudy (enricoanders):

c=69/50x+483/10

OpenStudy (enricoanders):

it has to be fractions :/

satellite73 (satellite73):

ok then what you got is right

OpenStudy (enricoanders):

okayy what about B

OpenStudy (enricoanders):

could you help with that?

satellite73 (satellite73):

where you see a T, put 80

OpenStudy (enricoanders):

ok

satellite73 (satellite73):

i would still use decimals, but whatever

OpenStudy (enricoanders):

so 552/5

OpenStudy (enricoanders):

its an online probram, it has to be fraction. it sucks

satellite73 (satellite73):

i get this http://www.wolframalpha.com/input/?i=69*80%2F50%2B483%2F10

OpenStudy (enricoanders):

WHAT im so lost

OpenStudy (enricoanders):

let me try it

OpenStudy (enricoanders):

its wrong

satellite73 (satellite73):

here is the line http://www.wolframalpha.com/input/?i=line+(-35,0),+(65,138)

satellite73 (satellite73):

and here it is for T = 80 http://www.wolframalpha.com/input/?i=483%2F10%2B(69+x)%2F50+for+x%3D80

satellite73 (satellite73):

did you use function notation?

OpenStudy (enricoanders):

what do you mean?

satellite73 (satellite73):

\[C(T)=...\]

OpenStudy (enricoanders):

yes :/

OpenStudy (enricoanders):

thank you tho

satellite73 (satellite73):

sorry not sure why it says you are wrong...